If x is an odd number and the sum of x and two consecutive odd numbers after x is 57, find the value of x.

Answer by oberobic(2304)

x is an odd number
x+2 is the next odd number
x+2+2 = x+4 is the next odd number
.
x +x+2 +x+4 = 57
.
3x +6 = 57
.
3x = 51
.
x = 17
x +2 = 19
x +4 = 21
.
17 +19 +21 = 57
.
Answer: The three consecutive odd numbers are 17, 19 and 21.
.
Done.